Let z= 4ex lny ,x= ln(ucosv), and y=usinv.  Calculate 𝜕z/𝜕u and 𝜕z/𝜕v first using the chain rule, and then by writing z(u,v) = z(x(u,v),y(u,v)) and taking partial derivatives directly. Show all steps each way


1
Expert's answer
2020-10-05T18:16:48-0400

"z=4e^x\ln y, x=\ln(u\cos v), y=u\sin v."


Calculate "\frac{\partial z}{\partial u}" and "\frac{\partial z}{\partial v}" using the chain rule:


"\frac{\partial z}{\partial u}=\frac{\partial z}{\partial x}\frac{\partial x}{\partial u}+\frac{\partial z}{\partial y}\frac{\partial y}{\partial u}"


"\frac{\partial z}{\partial u}= 4e^x\ln y\frac{\cos v}{u\cos v}+4e^x\frac{1}{y}\sin v= 4e^{\ln(u\cos v)}\ln(u\sin v)\frac{1}{u}+4e^{\ln(u\cos v)}\frac{1}{u\sin v}\sin v = 4u\cos v\ln(u\sin v)\frac{1}{u}+4u\cos v\frac{1}{u}=
4\cos v\ln(u\sin v)+4\cos v"


"\frac{\partial z}{\partial v}=\frac{\partial z}{\partial x}\frac{\partial x}{\partial v}+\frac{\partial z}{\partial y}\frac{\partial y}{\partial v}"


"\frac{\partial z}{\partial v}= 4e^x\ln y\frac{-u\sin v}{u\cos v}+4e^x\frac{1}{y}u\cos v=
 -4e^{\ln(u\cos v)}\ln(u\sin v)\frac{\sin v}{\cos v} +4e^{\ln(u\cos v)}\frac{1}{u\sin v}u\cos v=
 -4u\cos v\ln(u\sin v)\frac{\sin v}{\cos v} +4u\cos v\frac{1}{u\sin v}u\cos v=-4u\ln(u\sin v)\sin v+4u\frac{\cos^2 v}{\sin v}"

Write "z(u,v) = z(x(u,v),y(u,v))" and take partial derivatives directly:


"z=4e^{\ln(u\cos v)}\ln(u\sin v)=4u\cos v\ln(u\sin v)"


"\frac{\partial z}{\partial u}= 4\cos v\ln(u\sin v)+4u\cos v\frac{\sin v}{u\sin v}= 4\cos v\ln(u\sin v)+4\cos v"


"\frac{\partial z}{\partial v}=-4u\sin v\ln(u\sin v)+4u\cos v\frac{u\cos v}{u\sin v}=-4u\sin v\ln(u\sin v)+4u\frac{\cos^2 v}{\sin v}"